OpenStudy (anonymous):

find n. x^8 x x^n = x^20

jimthompson5910 (jim_thompson5910):

\[\Large x^8 \times x^n = x^{20}\] \[\Large x^{8+n}= x^{20}\] Since the bases are equal, the exponents are equal So 8 + n = 20 8 + n = 20 8 + n -8 = 20-8 n = 12

OpenStudy (anonymous):

thanks! could you help me with x^n x x^-3 =x^8

jimthompson5910 (jim_thompson5910):

sure thing \[\Large x^n \times x^{-3} = x^8\] \[\Large x^{n+(-3)} = x^8\] Since the bases are equal, the exponents are equal, so n+(-3) = 8 n+(-3) = 8 n-3 = 8 n-3+3 = 8 + 3 n = 11